I know these general prices: pulpwood is $7.00 per ton C/S is $12.00 per ton and sawtimber is $29.00 per ton. This ratio between prices should be maintained for the solution. I have determined that one acre of land has 20 tons of pulpwood 60 tons of C/S and 30 tons of sawtimber. I have sold the wood to someone who paid me a lump sum of $12,000..00. How much was paid for each unit of wood; pulpwood=?per ton, C/S =?per ton, and sawtimber = ? per ton. The answers should if possible result in the same ratio between prices as the beginning statement. I would also like to know how to solve this problem so that I can plug in the correct varriables.
Thanks,

Answer by stanbon(57377)

I know these general prices: pulpwood is $7.00 per ton C/S is $12.00 per ton and sawtimber is $29.00 per ton. This ratio between prices should be maintained for the solution. I have determined that one acre of land has 20 tons of pulpwood 60 tons of C/S and 30 tons of sawtimber. I have sold the wood to someone who paid me a lump sum of $12,000.00.
How much was paid for each unit of wood?
----------------
Ratio Equation: 7x + 12x + 29x = 12000
48x = 12000
x = 250
------------
7x = 175 (amount to spend for pulpwood)
12x= 3000 (amount to spend for C/S)
29x = 7250 (amount to spend for sawtimber)
---------------------------------------
Unit amount paid for pulpwood: 175/20 = $8.75 per ton
Unit amount paid for C/S: 3000/60 = $50 per ton
Unit amount paid for sawtimber: 250/30 = $8.33 per ton
==================
Cheers,
Stan H.
==================